The flow energy of 142 liters/min of a fluid passing a boundary to a system is 108.5 kJ/min. Determine the pressure at this point. 

p = ___ kPa

Answer in 2 decimal places


1
Expert's answer
2021-08-25T13:55:47-0400

Energy flow Q = 142 L/min

Boundary to system P= 108.5 kJ/min

"P= \frac{108.5 \times 10^3\;J}{60\;sec} \\

P= 1808.3 \;W \\

P=1.81 kW"

We know that,

"P= \frac{Work \; done \;(W)}{Time \; (t)} = \frac{Force \times distance}{Time} \\

P= Force \; (F) \times V \\

Velocity \; V = \frac{distance}{Time} \\

P=P_1 \times A \times V \\

Pressure \;(P_1) = \frac{F}{A} \\

P = P_1 \times Q \\

Energy \; flow \; Q = A \times V"

Putting values

"P_1 = \frac{P}{Q} = \frac{1.81 \;kW \times 60 \;sec}{0.142\;m^3/min} \\

= 764.78 \;kPa"

Answer: 764.78 kPa